# AuthPrint: Fingerprinting Generative Models Against Malicious Model Providers

### v1

@@ Abstract
Generative models are increasingly adopted in high-stakes domains, yet current deployments offer no mechanisms to verify the origin of model outputs. We address this gap by extending model fingerprinting techniques beyond the traditional collaborative setting to one where the model provider may act adversarially. To our knowledge, this is the first work to evaluate fingerprinting for provenance attribution under such a threat model. The methods rely on a trusted verifier that extracts secret fingerprints from the model’s output space, unknown to the provider, and trains a model to predict and verify them. Our empirical evaluation shows that our methods achieve near-zero FPR@95%TPR for instances of GAN and diffusion models, even when tested on small modifications to the original architecture and training data. Moreover, the methods remain robust against adversarial attacks that actively modify the outputs to bypass detection. Source codes are available at https://github.com/PSMLab/authprint.
Introduction
@@
Recent advances in generative AI have led to the widespread deployment of generative models across various domains, with providers of generative AI services increasingly monetizing their models by offering subscription-based access. However, this rapid adoption has raised serious concerns about the risks posed by these models, particularly in high-stakes and sensitive sectors, such as healthcare and defense, where erroneous model outputs can have disastrous consequences (Bommasani et al. 2021).
@@
In response, policymakers are introducing legal frameworks to regulate the use of AI and, in particular, the deployment of generative models. For instance, the European Union’s AI Act mandates independent, periodic audits for “high-risk” AI systems deployed in domains such as healthcare, education, employment, and critical infrastructure (European Parliament and Council of the European Union 2023). This requirement to pass or be certified by an audit raises a critical question: How can users verify that a given output indeed originated from the audited model? Even if a model passes an audit, there is currently no guarantee that providers will adhere to the audited model and not substitute it for a cheaper but potentially flawed version.
@@
Without verification mechanisms, audits may fail to enforce compliance. In particular, it becomes difficult to detect whether the model deployed after an audit matches the one audited. Similar challenges have been observed in other industries, where companies have deceived auditors by manipulating a product’s performance only to pass an audit, while deploying a product with detrimental effects on climate or public health (US EPA 2015). In the context of AI, service providers have similar intentions to reduce costs, possibly disregarding negative societal outcomes (Valdarrama 2023).
@@
This issue is exacerbated in current AI/ML systems, where access to the model is offered through opaque APIs that provide no verifiable information about which models are accessed. Users are left to blindly trust the model provider: for any single query, the model provider can plausibly deny having used a harmful or defective model, as little evidence exists to disprove their claims. The lack of accountability hinders oversight and diffuses responsibility across the ML supply chain (Veale 2023). Therefore, technical means to track models through their lifecycle—from certification to deployment—is critical, not only for law enforcement but also for maintaining trust in AI applications.
Prior work has proposed cryptographic commitment schemes based on zero-knowledge proofs to verify that an output is obtained by evaluating the committed model (Kang et al. 2023; Lee et al. 2024). Alternatively, trusted execution environments have been explored to prevent unauthorized modifications of a model and verify that computations originate from a specific model snapshot (Tramer and Boneh 2018). However, these technical approaches fall short of the requirements for deployment, as they either do not scale to the size of modern generative models, or require specialized hardware, limiting their potential adoption. Most prior work in this area targets small classifiers, overlooking the unique challenges posed by modern generative models.
@@
In this paper, we propose AuthPrint (Authentication via Fingerprinting), a novel framework for detecting the source model of generated images. AuthPrint relies on a trusted third party (e.g., an auditor) to train a model that learns to reconstruct secret fingerprints from images generated by the original audited model. During verification, the reconstructed fingerprint is compared against the true fingerprint to determine whether the image was produced by this authentic model. AuthPrint satisfies two key properties: (i) it reliably detects deviations in the output distribution from the original audited model; and (ii) it is robust against attacks attempting to forge images that pass verification.
Our main contributions are:
@@
  • We introduce AuthPrint, a black-box fingerprinting framework for attributing image outputs to a specific generative model, even under a malicious provider.
@@
  • We demonstrate its effectiveness across architectures and datasets, including StyleGAN2 (FFHQ, LSUN-Cat) and multiple Stable Diffusion versions.
@@
  • We show that AuthPrint remains robust against adaptive attackers with full access to the original model, reducing forgery success rates from 100% to zero.
@@
  • We provide a theoretical analysis of fingerprint recovery attacks, showing such attacks are infeasible for attackers with realistic constraints.
Background
@@
We consider a setting involving two entities: the model provider and the verifier. The model provider maintains a generative image model and monetizes it by offering remote access (e.g., via a web API). We assume the model has been trained and can be represented as a function mapping structured inputs to high-dimensional image outputs. More formally, we define an image generation model as a function $G_{\psi}:\Xi\times\mathcal{C}\to\mathbb{R}^{d}$, where $\psi$ denotes its parameters, $\Xi$ is the space of latent variables, $\mathcal{C}$ is the space of optional conditioning inputs, such as class labels or text prompts, and $\mathbb{R}^{d}$ is the output space corresponding to images of dimension $d=\text{channels}\times\text{height}\times\text{width}$. The parameters $\psi$ include all components required to fully specify the model for evaluation. For example, if $G_{\psi}$ is implemented as a neural network, $\psi$ would include its weights and architecture.
@@
Before deployment, the model provider must grant the verifier query access. We refer to this pre-deployment stage as the certification phase. During this phase, the verifier may issue a number of queries via the API, and the provider responds by evaluating $G_{\psi}$ on the verifier’s inputs and returning the resulting output images. We refer to the period after the model has been deployed as the verification phase. The central question we investigate is: Given any output $x\in\mathbb{R}^{d}$ at verification time, can the verifier determine whether it was produced by the same model used during certification?
@@
We envision real-world scenarios where this problem is relevant. Our running example is an AI audit, where the verifier is an auditor certifying the model’s compliance with AI regulations, such as fairness and privacy. Depending on how the regulations are implemented, the provider may grant the auditor query access, or extend the access under legal protections for intellectual property. Next, we detail the goals and capabilities of the model provider and verifier.
Threat Model
We consider a model provider aiming to maximize profit from its image generation service. The main threat we address is the provider acting maliciously and replacing the model with one of inferior quality but cheaper to evaluate, thereby reducing costs. The bogus model may produce images with flaws that are imperceptible to humans but may have a significant impact (e.g., quality of medical images in cancer diagnosis). In this paper, we focus on image quality deterioration, as it is often a direct consequence of evaluating a cheaper model. However, there are more subtle model flaws harder to detect on a single output, such as increased privacy risk or overrepresentation of certain demographics.
Verifier’s goal and capabilities. The verifier is only permitted black-box access to the model through a large number of queries at certification time. They may use this information to develop a detector $D$, such that $D(x)$ accurately detects whether or not $x$ was sampled from the certified model.
Adversary’s goal. The goal of the attack is to craft an output $\hat{x}\in\mathbb{R}^{d}$ in such a way that $\hat{x}$ deviates from the certified model’s output distribution, but $D$ incorrectly detects as drawn from it. Borrowing terminology from the ML watermarking literature, we call this attack a forgery attack, since if successful, the attacker deceives the verifier by making the sample appear to be drawn from the legitimate model.
Adversary’s capabilities. During certification, the provider must consistently respond using the same model. Note that the verifier’s goal is not to detect model defects, but to uncover a model swap. Moreover, the increased scrutiny during certification would be a deterrent to such misconduct. Second, the model provider will not spend more than the cost of sampling from the certified model, as that is against their goal of saving costs. Importantly, the adversary has full access to the certified model, which gives them a significant advantage in mounting forgery attacks.
@@
Model fingerprinting techniques
@@
Model fingerprinting refers to techniques that identify ML models based on distinctive patterns in their outputs. These methods are typically used to protect the intellectual property of model owners when models are stolen or leaked. By fingerprinting a model, unauthorized uses can be identified.
Most fingerprinting techniques fall into two categories: passive and active. Passive fingerprinting relies solely on detecting patterns in model outputs, while active fingerprinting modifies the model to make its outputs more identifiable.
We focus on passive techniques, as we assume the verifier only has query access during certification. While white-box access could enable active methods, these are vulnerable in adversarial settings: a provider with access to both original and fingerprinted models could conduct side-by-side comparisons to uncover the fingerprint or mount forgery attacks.
Passive fingerprinting. Prior work has proposed passive fingerprinting methods for model attribution. Marra et al. apply denoising filters to extract noise residuals from images, showing that the average residuals of a GAN correlate with its outputs and can serve as a fingerprint (Marra et al. 2019). Improving upon this, Yu et al. explore deep architectures to extract more discriminative features (Yu, Davis, and Fritz 2019). Corvi et al. identify patterns in the power spectra and autocorrelation of noise residuals of latent diffusion models (Corvi et al. 2023). Lastly, Song et al.’s geometric approach improves detection by approximating the projection of generated images onto the data manifold as vectors to their nearest counterparts in a real image dataset (Song, Khayatkhoei, and AbdAlmageed 2024).
Our approach differs from these works in two key ways. First, instead of a closed world of candidate models, we tackle attribution in an open-world: whether or not an image originates from the certified model. Second, we assume an adversarial model provider, as opposed to a trusted one. To our knowledge, we are the first to evaluate the susceptibility of these techniques to adaptive forgery attacks.
AuthPrint: Authenticity Auditing via Covert Fingerprinting
@@
Figure 1: Overview of the AuthPrint pipeline. In the certification phase, a verifier with black-box access to the claimed model selects secret pixel locations (channel-aware; simplified in figure for illustration) as a fingerprint and trains a private reconstructor to predict them from generated images. After certification, the provider deploys the model via its black-box API, while the verifier hosts a separate verification API with the reconstructor and secret indices. In the verification phase, a user submits an image to this API, which applies the reconstructor, compares predicted and actual values at the secret locations, computes a detection error, and returns the authenticity verdict.
@@
1: Sample secret index vector: $s\overset{\$}{\leftarrow}\mathcal{U}([d]^{l})$
@@
4:  Sample batch $\{(\xi_{i},c_{i})\}_{i=1}^{B}\overset{\$}{\leftarrow}P_{\mathrm{in}}$
@@
The core intuition behind AuthPrint lies in the fine-grained pixel-level dependencies learned by modern generative models (van den Oord, Kalchbrenner, and Kavukcuoglu 2016; Theis and Bethge 2015). A reconstructor trained to recover a secret subset of pixels from generated images performs well when the images are drawn from the distribution of training images (i.e., the certified model’s output distribution) but, if the provider substitutes the model with another generator, even slight shifts in these statistical dependencies result in prediction errors. AuthPrint exploits this out-of-distribution sensitivity by training the reconstructor on ample samples from the original model, enabling robust detection of distributional shifts without modifying the model or images. It can be directly applied to a wide range of models and learning algorithms, requires no retraining of the generator, is hardware-independent (unlike secure hardware-based approaches (Tramer and Boneh 2018)), and remains practical for high-capacity models where cryptographic approaches such as zero-knowledge proofs are infeasible.
The security of AuthPrint rests on the secrecy of the detector, namely the reconstructor parameters $\phi$ and the index vector $s$. The model provider is aware of AuthPrint’s algorithms and may query the detector during the verification phase, e.g., impersonating a user of the verification API, but the API only provides binary responses. Importantly, the model provider has no access to gradients, logits, or auxiliary information about the reconstructor. Our evaluation shows that this information asymmetry thwarts forgery and reverse-engineering attacks, ensuring the robustness of AuthPrint even when deployed in an adversarial setting.
Results
Detection Performance on Model Substitution
We first evaluate AuthPrint’s detection performance on model substitution for both GAN-based (StyleGAN2) and diffusion-based (SD) models (see Appendix A for experiment details). To test rejection of negative models, we report detection error as the FPR at 95% TPR, treating target model samples as positives. As an ablation, we vary the fingerprint length (i.e., number of secret pixels) to assess its effect on detection error, holding all other factors constant [^1] Variance across random seeds is $1\%$; we report a single run..
In the StyleGAN2 experiments, the target and negative models share the same architecture but differ in training configurations such as dataset size and data augmentation. This setup probes AuthPrint’s sensitivity to subtle changes in the training procedure that induces negligible image quality shifts. We evaluate on FFHQ and LSUN-Cat, with results shown in Figure 2. For SD, the target is SD 2.1, and negatives are earlier versions (SD 1.5 to 1.1), introducing progressively larger distributional shifts due to architectural and dataset differences. Results are shown in Figure 3.
Figure 2: AuthPrint detection performance on model substitution for pre-trained StyleGAN2 models (left: FFHQ, right: LSUN-Cat) across varying fingerprint lengths. Target and negative models share the architecture but differ in training data size or augmentation, as detailed in the figure.
AuthPrint consistently achieves low detection error across both model families, validating its robustness in distinguishing target from non-target outputs. Detection performance peaks at fingerprint lengths of 16 and 32 for StyleGAN2 and 64 for SD. Short fingerprints yield insufficient signal, while overly long ones introduce redundancy that limits reconstructor accuracy under fixed capacity.
Figure 3: AuthPrint detection performance on model substitution for SD models across varying fingerprint lengths. The target is SD 2.1; negative models are earlier versions (SD 1.5, 1.4, 1.3, 1.2, 1.1) with lower image qualities. The Japanese café prompt is used throughout.

### v2

@@ Abstract
Generative models are increasingly adopted in high-stakes domains, yet current deployments offer no mechanisms to verify whether a given output truly originates from the certified model. We address this gap by extending model fingerprinting techniques beyond the traditional collaborative setting to one where the model provider itself may act adversarially, replacing the certified model with a cheaper or lower-quality substitute. To our knowledge, this is the first work to study fingerprinting for provenance attribution under such a threat model. Our approach introduces a trusted verifier that, during a certification phase, extracts hidden fingerprints from the authentic model’s output space and trains a detector to recognize them. During verification, this detector can determine whether new outputs are consistent with the certified model, without requiring specialized hardware or model modifications. In extensive experiments, our methods achieve near-zero FPR@95%TPR on both GANs and diffusion models, and remain effective even against subtle architectural or training changes. Furthermore, the approach is robust to adaptive adversaries that actively manipulate outputs in an attempt to evade detection.
I Introduction
@@
Recent advances in generative AI have led to the widespread deployment of generative models across various domains, with providers of generative AI services increasingly monetizing their models by offering subscription-based access. However, this rapid adoption has raised serious concerns about the risks posed by these models, particularly in safety-critical domains, such as healthcare and defense, where erroneous model outputs can have disastrous consequences [1].
@@
In response, policymakers are introducing legal frameworks to regulate the use of AI and, in particular, the deployment of generative models. For instance, the European Union’s AI Act mandates independent, periodic audits for “high-risk” AI systems deployed in domains such as healthcare, education, employment, and critical infrastructure [2]. This requirement to pass or be certified by an audit raises a critical question: How can users verify that a given output indeed originated from the audited model? Even if a model passes an audit, there is currently no guarantee that providers will adhere to the audited model and not substitute it for a cheaper but potentially flawed version.
@@
Without verification mechanisms, audits may fail to enforce compliance. In particular, it becomes difficult to detect whether the model deployed after an audit matches the one that had been audited. Similar challenges have been observed in other sectors, where companies have deceived auditors by manipulating a product’s performance only to pass an audit, while deploying a product with detrimental effects on climate and public health [3]. In the context of AI, service providers have similar intentions to reduce costs, possibly disregarding harmful societal outcomes [4].
@@
This issue is exacerbated in current AI/ML systems, where access to the model is offered through opaque APIs that provide no verifiable information about which models are accessed. Users are left to blindly trust the model provider: for any single query, the model provider can always plausibly deny having used a harmful or defective model, as users have no access to evidence that could disprove these claims. This lack of accountability hinders oversight and diffuses responsibility across the ML supply chain [5, 6]. Therefore, technical means to track models through their lifecycle—from certification to deployment—is critical, not only for law enforcement but also for maintaining trust in AI applications.
Several cryptographic methods have been proposed in the literature to verify that outputs originate from a specific model. While early cryptographic verification protocols were highly interactive and thus not applicable on single queries [7], recent breakthroughs in the field of zero-knowledge have enabled proofs that can verify that a previously committed model is evaluated on a given query [8]. Recent zk-ML systems compile small networks into arithmetic circuits to enable zk-SNARKs [9, 10, 11, 12, 13], but the overheads of these systems—in terms of the added latency of proof verifications— still limit their practical feasibility, particularly in domains where fast responses are critical.
Another line of work binds models to trusted execution environments (TEEs), with outputs verified via remote attestation [14]. Yet GPU-based TEEs remain immature, and hybrid schemes that delegate computations to untrusted GPUs require weight quantization, adding latency and reducing accuracy. Overall, both cryptographic and TEE-based solutions face scalability, efficiency, and hardware barriers, making them impractical in many deployments of large generative models.
@@
In this paper, we propose AuthPrint (Authentication via Fingerprinting), a novel framework for detecting the source model of generated images. AuthPrint relies on a trusted third party (e.g., an auditor) to train a model that learns to reconstruct secret fingerprints from images generated by the original audited model. During verification, the reconstructed fingerprint is compared against the true fingerprint to determine whether the image was produced by this authentic model.
Crucially, our empirical evaluation shows that AuthPrint satisfies two essential properties: (i) it reliably detects deviations in output distributions from the original audited model, and (ii) it remains robust against attacks aimed at forging images that pass verification. We show that AuthPrint achieves consistently low detection error on both StyleGAN2 and Stable Diffusion models, even under model substitutions with minor training modifications, model compression, and adaptive adversarial attacks based on evasion and fingerprint recovery. These results position AuthPrint as an effective method for authenticating image generation models across diverse architectures and adversarial settings.
To sum up, our main contributions are:
@@
  • We introduce AuthPrint, a black-box covert fingerprinting framework for attributing image outputs to a specific generative model, even in the presence of a malicious model provider.
@@
  • We demonstrate its effectiveness across diverse architectures and datasets, including StyleGAN2 (trained on FFHQ and LSUN-Cat) and multiple versions of Stable Diffusion.
@@
  • We show that AuthPrint is robust to adaptive adversaries with full access to the original audited model, reducing forgery success rates from 100% (in baseline methods) to zero.
@@
  • We provide a theoretical analysis of fingerprint recovery attacks, establishing their infeasibility under realistic adversarial constraints.
II Background
@@ II-A Problem Statement
We consider a setting involving two entities: the model provider and the verifier. The model provider maintains a generative image model and monetizes it by offering remote access (e.g., via a web API). We assume the model has been trained and can be represented as a function mapping structured inputs to high-dimensional image outputs. More formally, we define an image generation model as a function $G_{\psi}:\Xi\times\mathcal{C}\to\mathbb{R}^{d}$, where $\psi$ denotes its parameters, $\Xi$ is the space of latent variables, $\mathcal{C}$ is the space of conditioning inputs, such as class labels or text prompts, and $\mathbb{R}^{d}$ is the output space corresponding to images of dimension $d=\text{channels}\times\text{height}\times\text{width}$. The parameters $\psi$ include everything required to fully specify the model for inference. For example, if $G_{\psi}$ is implemented as a neural network, $\psi$ would include its weights and architecture.
@@
Before deployment, the model provider must grant the verifier query access. We refer to this pre-deployment stage as the certification phase. During this phase, the verifier may issue a number of queries via the API, and the provider responds by evaluating $G_{\psi}$ on the verifier’s inputs and returning the resulting output images. We refer to the period after the model has been deployed as the verification phase. The central question we investigate is: Given any output $x\in\mathbb{R}^{d}$ at verification time, can the verifier determine whether it was produced by the model that was certified?
@@
We envision real-world scenarios where this problem is relevant. Our running example is an AI audit, where the verifier is an auditor certifying the model’s compliance with AI regulations, such as the need to meet appropriate fairness and privacy requirements for the specific application. Depending on how the regulations are implemented, the provider may grant the auditor query access, or provide full access to the model under legal non-disclosure agreements. Next, we detail the goals and capabilities of the model provider and verifier.
II-B Threat Model
We consider a model provider aiming to maximize profit from its image generation service. The main threat we address is post-certification model substitution: the provider may deploy, behind an opaque API, a different generator of inferior quality but cheaper to evaluate. Such a bogus model can introduce flaws that are imperceptible in single images yet material for downstream use (e.g., degraded medical image fidelity). In this work we focus on distributional shifts that manifest as quality deterioration, while noting that other shifts (e.g., privacy leakage or demographic skews) are also plausible.
Certification-phase authenticity assumption. Our setting targets deployments subject to oversight (e.g., regulatory or contractual audits) in which the provider grants a verifier controlled query access before deployment. During this certification phase, the verifier interacts with the provider’s model instance to build a detector $D$. We assume the instance exposed at certification is the authentic model intended for deployment. This mirrors standard audit practice: access occurs under legal or commercial obligations that penalize substitution at certification time. The assumption is operationally realistic and can be reinforced with lightweight safeguards during certification, such as session binding or simple challenge–response checks, though these are not required at verification. Our contribution then addresses the post-certification risk: once the model is deployed behind an opaque API, $D$ must enable users to verify that outputs still come from the certified distribution.
Verifier’s goal and capabilities. The verifier that we consider has only black-box access during certification but can issue many queries to the authentic model to train a detector $D$ that, given an image $x$, decides whether $x$ was sampled from the certified model’s output distribution. The verifier does not alter the provider’s model and does not rely on specialized hardware. After deployment, the verifier exposes $D$ as a separate black-box verification API that returns only a binary output about the authenticity of the queried model.
Adversary’s goal. After certification, the provider may replace the authentic generator with a cheaper substitute or a modified variant and still have outputs accepted by $D$. Concretely, the adversary aims to produce $\hat{x}\in\mathbb{R}^{d}$ whose distribution deviates from that of the certified model while $D$ incorrectly accepts it as authentic (a forgery attack).
Adversary’s knowledge and capabilities. During certification, the provider must answer queries using a single fixed model instance; attempting a swap during the certification phase can be deterred by oversight and additional safeguards. Post-certification, the provider has full white-box access to the authentic model. However, the provider has no access to the internals of the verifier’s private detector and can only see the binary accept/reject outputs of their queries—which are likely to be rate-limited by the verifier.
We assume the adversary is not constrained in the architecture or training choices they can make to develop the substitute model, except that the cost of evaluating the substitute model should not exceed that of the authentic model, as the motivation for substitution is cost saving. This rules out attacks that are systematically more expensive than sampling from the certified generator.
In summary, we assume authentic access only during a controlled, pre-deployment certification window—an assumption aligned with real audit workflows—and we design $D$ to detect subsequent model substitution when users can see only the provider’s opaque API outputs.
@@ II-B Threat Model
II-C Model Fingerprinting Techniques
@@
Model fingerprinting refers to methods that identify machine learning models by exploiting distinctive statistical patterns in their outputs. These methods are typically motivated by intellectual property protection: when a model is stolen, leaked, or re-deployed without authorization, its fingerprint can be used to trace the source model.
Fingerprinting techniques are commonly divided into two categories: active and passive. Active fingerprinting modifies the model itself—for example, by embedding secret perturbations in the weights—so that its outputs carry an identifiable signal. While effective in cooperative or white-box settings, active methods are fragile when the provider may behave adversarially: access to both the fingerprinted and unmodified models allows side-by-side comparisons to expose or strip the fingerprint.
By contrast, passive fingerprinting does not alter the model, but instead relies on statistical regularities that naturally emerge in its outputs. This makes passive approaches better suited for our setting, where the verifier only has black-box access to the model during certification.
Passive fingerprinting for model attribution. A growing body of work has explored passive model fingerprinting techniques for attributing generated images to their source model. Yu et al. demonstrate that noise residuals of generated images contain generator-specific traces, and propose deep architectures that exploit these patterns to identify which of several GANs produced a given image [15]. Dzanic et al. show that generative models leave distinctive patterns in the Fourier spectrum—particularly in the high frequency band—that can serve as reliable fingerprints to distinguish different models [16]. Girish et al. extend attribution to scenarios where new generators may emerge post-training, by combining out-of-distribution detection and clustering to discover and assign samples from previously unseen models [17]. More recently, Song et al. take a geometric approach in which generated samples are projected onto the manifold of real images, with the resulting offsets serving as model-specific fingerprints [18]. Follow-up work generalizes this approach to Riemannian geometry, yielding more faithful representations of data structure and improved attribution across diverse datasets and architectures [19].
These studies typically report model attribution accuracies in the range of 70–99% with a pool of between 5 to 20 candidate models. Although they are evaluated in small worlds of models, the techniques show promise for reliable model identification, aligning with the objectives of our work. However, these techniques are not suitable to our setting: we consider an adversarial setting where an untrusted provider may substitute the certified model after deployment to reduce costs or evade oversight. In this case, the adversary has full access to both the audited and substitute models, and the fingerprinting techniques. Therefore, if these fingerprinting methods were used, the adversary could easily replicate or mimic the fingerprints, thereby bypassing detection.
To our knowledge, no existing passive fingerprinting approaches have been designed to address such an adversarial setting, nor have they been evaluated under adaptive forgery attacks. Furthermore, prior evaluations often compare across heterogeneous model families (e.g., different GAN architectures or distinct generator types), whereas our setting requires the techniques to be robust against far subtler modifications, such as variants trained with slightly different configurations or models subjected to compression techniques. These gaps motivate the approach we introduce in the next section.
III AuthPrint: Authenticity Auditing via Covert Model Fingerprinting
@@
Figure 1: Overview of the AuthPrint pipeline. In the certification phase, a verifier with black-box access to the claimed model selects secret pixel locations (image channels are simplified for illustration) as a fingerprint and trains a private reconstructor to predict them from generated images. After certification, the provider deploys the model via its black-box API, while the verifier hosts a separate verification API with the reconstructor and secret indices. In the verification phase, a user submits an image to this API, which applies the reconstructor, compares predicted and actual values at the secret locations, computes a detection error, and returns the authenticity verdict.
@@
1: Sample secret index vector: $s\overset{\mathdollar}{\leftarrow}\mathcal{U}([d]^{l})$
@@
4:  Sample batch $\{(\xi_{i},c_{i})\}_{i=1}^{B}\overset{\mathdollar}{\leftarrow}P_{\mathrm{in}}$
@@
The core intuition behind AuthPrint rests on two factors. First, modern generative models exhibit fine-grained pixel-level dependencies [20, 21], which enable a reconstructor (a deep neural network) to accurately recover a secret subset of pixels from the full image. Second, we exploit the out-of-distribution (OOD) sensitivity of deep neural networks: the reconstructor performs reliably only when images are sampled from the certified model’s distribution but fails under substitution, as even subtle shifts in conditional dependencies introduce prediction errors. This approach to detecting distribution shifts provides a principled mechanism for determining whether outputs originate from the claimed model.
To instantiate the fingerprint, we sample the index vector $s$ uniformly at random. Random selection ensures that the fingerprint probes pixel dependencies without bias toward specific regions or channels and avoids predictable structures that an adversary might exploit. It also maximizes adversarial uncertainty: if $s$ were fixed or exhibited any discernible pattern, a provider could evade detection by training a substitute reconstructor for gradient-based attacks or by strategically adjusting the chosen pixels. By keeping $s$ both random and secret, any successful forgery must implicitly preserve the full joint distribution of pixels, since the adversary cannot anticipate which subset will be verified.
The security of AuthPrint therefore rests on the secrecy of the detector, namely the reconstructor parameters $\phi$ and the index vector $s$. Although the model provider knows the verification protocol and may interact with the detector, they only observe binary outcomes. Without access to logits, gradients, or the indices themselves, they cannot train surrogates or mount gradient-based evasion attacks. As we show in later sections, recovering $s$ reduces to a threshold group-testing problem to identify the fingerprint set, followed by factorial search over all permutations to recover its order—an infeasible task at practical fingerprint lengths. This information asymmetry, combined with the unpredictability of random index selection, underpins AuthPrint’s robustness against forgery attacks with either evasion or reverse-engineering attempts.
IV Experimental Setup
We evaluate AuthPrint on two representative families of image generators: GAN-based and diffusion-based models, instantiated by StyleGAN2 and Stable Diffusion (SD), respectively. Our evaluation spans a range of adversarial behaviors, including model substitution, model modification, and adaptive attack scenarios. This section provides details on the models, datasets, training configurations, and attack setups.
Generative Models.StyleGAN2: We use models from the official StyleGAN2-ADA repository by NVLabs [^1] https://github.com/NVlabs/stylegan2-ada-pytorch, trained on FFHQ and LSUN-Cat, each producing $256\times 256$ images. For each dataset, we designate one target model and treat variants trained under different configurations as negative models:
  • FFHQ:
    • Target model: trained on 70k images with ADA (Adaptive Discriminator Augmentation) [22].
    • Negative models: trained on
        1k images with ADA
        30k images with ADA
        70k images with BCR (Balanced Consistency Regularization) [23]
        70k images with no augmentation
  • LSUN-Cat:
    • Target model: trained on 100k images with ADA.
    • Negative models: trained on
        1k images with ADA
        30k images with ADA
        100k images with BCR
        100k images with no augmentation
ADA is an adaptive augmentation strategy applied during discriminator training [22], while BCR is a regularization method that enforces consistency across augmented samples [23]. In both datasets, negative models differ in training set size or augmentation strategy, resulting in subtle but detectable distribution shifts.
Stable Diffusion: We use SD 2.1 as the target model and earlier versions (SD 1.5–1.1) as negatives, obtained from HuggingFace [^2] https://huggingface.co. Unless otherwise noted, these models generate $1024\times 1024$ images.
Fingerprint Reconstructor. The reconstructor is a CNN trained to predict pixel values at a secret subset of locations given the full generated image. The architecture of the CNN follows a conventional design: a convolutional encoder with stride-2 downsampling, $4\times 4$ kernels, batch normalization, and LeakyReLU activations, followed by global average pooling and an MLP head. For StyleGAN2 experiments, we use a 32M-parameter reconstructor, which is sufficient given the relatively low diversity and resolution of StyleGAN2 outputs. In contrast, SD experiments employ a larger 674M-parameter model to better capture the higher complexity and semantic richness of text-to-image generations. To study the effect of scaling, we additionally evaluate reconstructors with 32M, 187M, and 674M parameters on SD.
Training Data. For StyleGAN2, training data are generated by sampling latent vectors $z\sim\mathcal{N}(0,I)$. For SD, we default to a fixed, high-specificity prompt—referred to as the Japanese café prompt—chosen for its repeatability and richness of visual detail. This prompt was generated using OpenAI’s ChatGPT-4o model [24]: